# LLaVA-RadZ: Can Multimodal Large Language Models Effectively Tackle Zero-shot Radiology Recognition?

### v1

@@ Abstract
Recently, multimodal large models (MLLMs) have demonstrated exceptional capabilities in visual understanding and reasoning across various vision-language tasks. However, MLLMs usually perform poorly in zero-shot medical disease recognition, as they do not fully exploit the captured features and available medical knowledge. To address this challenge, we propose LLaVA-RadZ, a simple yet effective framework for zero-shot medical disease recognition. Specifically, we design an end-to-end training strategy, termed Decoding-Side Feature Alignment Training (DFAT) to take advantage of the characteristics of the MLLM decoder architecture and incorporate modality-specific tokens tailored for different modalities, which effectively utilizes image and text representations and facilitates robust cross-modal alignment. Additionally, we introduce a Domain Knowledge Anchoring Module (DKAM) to exploit the intrinsic medical knowledge of large models, which mitigates the category semantic gap in image-text alignment. DKAM improves category-level alignment, allowing for accurate disease recognition. Extensive experiments on multiple benchmarks demonstrate that our LLaVA-RadZ significantly outperforms traditional MLLMs in zero-shot disease recognition and exhibits the state-of-the-art performance compared to the well-established and highly-optimized CLIP-based approaches.
@@ 1 Introduction
With the rapid advancement of deep learning technologies, an increasing number of studies have focused on their applications in medical disease diagnosis, yielding remarkable results [1, 2, 3, 4]. However, these approaches typically rely on high-quality annotations provided by clinical experts. Unlike natural image datasets, annotating medical images is both costly and time-consuming. To address this challenge, recent research has explored methods based on paired medical images and textual reports, leveraging contrastive learning techniques. By minimizing the distance between paired samples while maximizing the distance between unpaired ones, these approaches enable zero-shot disease recognition, thereby reducing reliance on extensive medical data annotation to a certain extent.
@@
In our in-depth investigation of advanced zero-shot disease recognition methods in the medical domain, we examined several representative models, including CARZero [5], MedKLIP [6], KAD [7], and MAVL [8]. While these methods have achieved significant performance improvements, they invariably incorporate external models or expert domain knowledge to some extent, rather than fully leveraging the models’ intrinsic understanding capabilities. Recently, multimodal large language models (MLLMs) [9, 10, 11, 12, 13, 14] have demonstrated remarkable capabilities across various user-oriented vision-language tasks, such as image comprehension and reasoning, offering new possibilities for zero-shot disease recognition in medical applications. Among these, LLaVA-Med [15] has exhibited exceptional domain-specific medical knowledge in dialogue-based tasks, indicating that it possesses a certain degree of medical expertise. Therefore, building upon LLaVA-Med, we further explore its potential in zero-shot medical classification.
@@
However, recent study [16] found that MLLMs ( i.e., LLaVA [11]) performed significantly worse than CLIP [17] on standard image classification tasks. To further validate this observation, we conducted zero-shot classification experiments using multiple MLLMs on five medical imaging datasets. The experimental results are consistent with previous findings, confirming that MLLMs exhibit suboptimal performance in image classification, particularly when dealing with complex medical images. To enhance the generalization capability of MLLMs in radiology disease recognition tasks, we employed a fine-tuning strategy and performed supervised fine-tuning on the MIMIC-CXR dataset [18]. Additionally, inspired by the work of Zhang et al. [16], we incorporated a series of optimizations. While these improvements yielded performance gains, the results remained inferior compared to traditional CLIP-based models. This phenomenon raises a critical question: Can MLLMs effectively perform zero-shot disease recognition?
@@
To answer the aforementioned question, we visualize the feature distributions of LLaVA and MAVL [8] on the RSNA [19] dataset, as illustrated in Fig. 1. The results indicate that LLaVA exhibits strong feature extraction capabilities, comparable to the well-established MAVL in the domain. However, in the disease recognition task, MAVL significantly outperforms LLaVA. We hypothesize that this performance gap arises because MLLMs fail to fully utilize the extracted features for effective disease identification. To address this limitation, we propose a simple yet effective LLaVA-RadZ framework for zero-shot disease recognition. Firstly, we introduce a new end-to-end training strategy, Decoding-Side Feature Alignment Training (DFAT). Specifically, we introduce special tokens for both image and text modalities and leverage the autoregressive generation capability of the decoder architecture to extract global representations of images and texts. Additionally, we incorporate a cross-modal contrastive loss to optimize the model’s ability to learn discriminative features. Furthermore, to mitigate the semantic category gap encountered during fine-grained alignment between medical images and textual reports, we design a Domain Knowledge Anchoring Module (DKAM). DKAM utilizes the model’s intrinsic medical knowledge to extract the semantic information underlying disease categories, constructing disease description vectors that serve as an intermediary bridge to facilitate the alignment between medical images and textual reports, thereby establishing a stable trimodal relationship. To further enhance the correlation among medical images, textual reports, and disease categories, a category knowledge-guided loss strengthens the association between similar images and corresponding textual reports.
@@
We propose a novel end-to-end training strategy, Decoding-Side Feature Alignment Training (DFAT), and incorporate a cross-modal contrastive loss to optimize the model’s ability to achieve effective alignment between visual and textual features.
We design a Domain Knowledge Anchoring Module (DKAM) to leverage MLLM’s intrinsic medical knowledge and effectively mitigate semantic gap in image-text alignment, thereby enhancing category-level alignment.
@@
| Dataset | CheXpert [20] | ChestXray-14 [21] | COVIDx CXR-2 [22] | RSNA Pneumonia [19] | SIIM-ACR [23] | |||||||||||
| MLLM | LLaVA-1.5 (7B) [11] | - | 7.50 | 8.28 | - | 3.33 | 6.92 | - | 53.14 | 50.28 | - | 40.53 | 55.34 | - | 23.66 | 50.36 |
| LLavA-Med (7B) [15] | - | 6.87 | 8.94 | - | 8.02 | 6.78 | - | 34.90 | 50.03 | - | 18.58 | 50.00 | - | 21.91 | 49.90 | |
| Qwen2.5-Max [12] | - | 32.23 | 67.97 | - | 19.04 | 76.19 | - | 75.91 | 76.81 | - | 43.58 | 43.59 | - | 64.70 | 72.57 | |
| Gemini-Pro [10] | - | 35.01 | 76.08 | - | 14.16 | 77.78 | - | 62.84 | 62.90 | - | 44.23 | 51.43 | - | 61.43 | 72.03 | |
| GPT-4o [9] | - | 45.85 | 81.14 | - | 19.85 | 81.55 | - | 50.93 | 77.08 | - | 54.20 | 65.33 | - | 64.57 | 72.11 | |
| LLavA-Med-7B + CoT [16] | - | 8.90 | 26.23 | - | 8.33 | 20.46 | - | 27.12 | 26.55 | - | 49.59 | 43.80 | - | 54.06 | 51.07 | |
| LLavA-Med-7B + Inference [16] | 71.00 | 44.85 | 75.45 | 64.30 | 21.73 | 70.86 | 71.07 | 69.84 | 60.39 | 77.51 | 69.85 | 72.90 | 71.25 | 68.26 | 71.27 | |
| Ours | LLaVA-RadZ | 75.63 | 48.41 | 81.77 | 70.75 | 27.86 | 84.57 | 81.76 | 76.65 | 74.41 | 85.94 | 75.84 | 81.35 | 86.32 | 76.89 | 79.28 |
2 Related Work
Multi-modal Large Language Models. Inspired by the exceptional reasoning capabilities of large language models (LLMs), researchers are actively exploring ways to extend these abilities to the visual domain, driving advancements in multimodal LLMs. With the release of GPT-4 (Vision) [9] and Gemini [10], these models have demonstrated remarkable multimodal understanding and generation capabilities, further fueling research in this field.
To bridge the gap between vision encoders and LLMs, BLIP-2 [24] introduces a Q-Former that transforms image features into a format compatible with LLMs, enabling seamless integration with text embeddings. LLaVA [11] and MiniGPT-4 [25] further enhance generalization and task performance by leveraging large-scale multimodal pretraining, followed by instruction tuning for specific applications. In the medical domain, LLMs have shown immense potential for advancing research and practical applications. Med-Flamingo [26] extends Flamingo to the medical field by pretraining on multimodal knowledge sources spanning multiple medical disciplines. LLaVA-Med [15] refines image-text pairs from PMC-15M [27] and trains a biomedical-specialized MLLM using a limited dataset, building upon the pre-trained parameters of LLaVA. Similarly, Med-PaLM [28] fine-tunes PaLM [29] using domain-specific medical instructions, demonstrating strong performance under human evaluation frameworks. Other notable models, such as Chat-Doctor [30] and Med-Alpaca [31], have been tailored for medical question-answering and dialogue applications.
Despite the significant progress of MLLMs, several challenges remain [32, 33, 16, 34]. Recent studies [16, 34] highlight the suboptimal performance of MLLMs in image classification, particularly in fine-grained category recognition. We find that this issue is especially pronounced in the medical domain, where precise classification is crucial for medical applications. To address these shortcomings, we are refining traditional MLLM training paradigms to enhance classification performance and improve fine-grained category comprehension.
Prompt Engineering. Prompting enhances the ability of pre-trained large language models (LLMs) to understand tasks by incorporating language instructions into the input text [35, 36, 37, 38]. Recently, prompt-based techniques have also been applied to vision-language models to improve performance. In medical vision-language models (VLMs), GloRIA [39] generates a set of textual prompts to describe potential subtypes, severity levels, and anatomical locations for each disease category. MedKLIP [6] enhances model performance by retrieving descriptions of medical entities from the UMLS knowledge base [40]. CARZero [5] introduces a prompt-alignment strategy based on LLMs, integrating prompt templates into the training dataset to ensure alignment during both training and inference. MAVL [8] uses visual descriptions of pathological features to guide the model in effectively detecting diseases in medical images.
Although these approaches have successfully improved model performance through prompt-based strategies, they all rely on external models or expert knowledge, without fully leveraging the model’s intrinsic understanding capabilities. Fortunately, recent research on LLaVA-Med [15] has demonstrated remarkable domain-specific conversational abilities, proving that it possesses a certain level of medical knowledge. Building upon LLaVA-Med [15], we further explore the feasibility of utilizing the model’s inherent comprehension to enhance zero-shot medical classification performance.
@@ 3.1 Can Med-LLMs Be Good Medical Classifiers?
Previous studies have explored the classification capabilities of multimodal large language models (MLLMs), revealing that their performance on image classification tasks is often limited. For example, Zhang et al. [16] investigates the performance differences in classification between MLLMs and CLIP, focusing on factors such as inference strategies, training approaches, and datasets. Inspired by this work, we extend the exploration to zero-shot tasks in the medical domain. Unlike natural images and text, the relationship between medical images and reports is more complex. We seek to investigate whether large medical models, leveraging domain-specific knowledge, can achieve superior performance on medical zero-shot tasks.
We first evaluated two open-source MLLMs, i.e., LLaVA-1.5 [11] and LLaVA-Med [15], on five medical datasets in a zero-shot classification setting. The evaluation followed a general large-model classification approach, where the model selects the correct category from a set of candidate options. As shown in Tab. 1, these models demonstrated limited performance in disease classification tasks and failed to accurately identify various medical conditions. Given the potential knowledge limitations of these models, we further assessed the performance of more powerful proprietary MLLMs ( i.e., Qwen2.5-Max [12], Gemini-Pro [10], and GPT-4o [9]) on zero-shot medical disease recognition tasks. As shown in Tab. 1, these models exhibited superior classification capabilities. However, they still lagged behind the state-of-the-art domain-specific methods in medical classification.
To enhance the generalization ability of MLLMs in radiology disease identification, we conducted Supervised Fine-Tuning (SFT) on LLaVA-1.5 [11] and LLaVA-Med [15] using the publicly available MIMIC-CXR dataset [18]. Surprisingly, the fine-tuned models did not achieve consistent performance improvements across the five datasets. In some cases, their classification performance even deteriorated. Further analysis of the model outputs revealed that MLLMs did not always focus on disease-specific information in radiology reports. Instead, they tended to overlearn the textual structures and linguistic patterns of the reports, which limited their classification capability. To mitigate this issue, we incorporated the Chain-of-Thought (CoT) prompting strategy and adjusted the model’s reasoning approach, inspired by the methodology of Zhang et al. [16], to optimize the model’s decision-making process. This approach led to moderate improvements in classification performance on medical datasets. Although the models have not yet reached optimal performance, the results suggest that MLLMs still hold significant potential for zero-shot medical disease recognition.
@@ 3.2 Motivation
To address this limitation, we propose the LLaVA-RadZ framework, introducing a novel end-to-end training strategy, Decoding-Side Feature Alignment Training (DFAT). This approach leverages the unique properties of the MLLM decoder architecture while incorporating modality-specific special tokens to facilitate effective interaction between medical images and textual features, ultimately achieving more robust cross-modal alignment. As illustrated in Fig. 2, we compare the feature distribution of our model with MAVL [8], the current state-of-the-art method, on the RSNA [19] dataset. The results clearly demonstrate that our model achieves better clustering of intra-class samples while enhancing inter-class separation, validating the effectiveness of our approach. Furthermore, we introduce the Domain Knowledge Anchor Module (DKAM), which harnesses the intrinsic medical knowledge of LLMs to bridge the semantic gap between images and text, enabling more precise disease classification.
@@ 3.3.2 Domain Knowledge Anchor Module
| Dataset | ChestXray-14 [21] | COVIDx CXR-2 [22] | RSNA Pneumonia [19] | SIIM-ACR [23] | ||||||||
| Method | AUC | F1 | ACC | AUC | F1 | ACC | AUC | F1 | ACC | AUC | F1 | ACC |
| ConVIRT [41] | 53.15 | 12.38 | 57.88 | 62.78 | 71.23 | 63.84 | 79.21 | 55.67 | 75.08 | 64.25 | 42.87 | 53.42 |
| GLoRIA [39] | 55.92 | 14.20 | 59.47 | 64.52 | 70.78 | 60.21 | 70.37 | 48.19 | 70.54 | 54.71 | 40.39 | 47.15 |
| BioViL [42] | 57.82 | 15.64 | 61.33 | 61.40 | 70.92 | 58.20 | 84.12 | 54.59 | 74.43 | 70.28 | 46.45 | 68.22 |
| CheXzero [43] | 66.99 | 21.99 | 65.38 | 73.13 | 76.13 | 71.45 | 85.13 | 61.49 | 78.34 | 84.60 | 65.97 | 77.34 |
| MedKLIP [6] | 72.33 | 24.18 | 79.40 | 76.28 | 76.54 | 71.96 | 86.57 | 63.28 | 79.97 | 89.79 | 72.73 | 83.99 |
| MAVL [8] | 73.50 | 26.25 | 82.77 | 83.86 | 81.73 | 78.07 | 86.91 | 63.41 | 82.42 | 92.04 | 77.95 | 87.14 |
| Ours | 70.75 | 27.86 | 84.57 | 81.76 | 76.65 | 74.41 | 85.94 | 75.84 | 81.35 | 86.32 | 76.89 | 79.28 |
| Dataset | RSNA Pneumonia [19] | Pneumothorax [23] | COVIDx CXR-2 [22] | ||||||
| Data portion | 1% | 10% | 100% | 1% | 10% | 100% | 1% | 10% | 100% |
@@
| ConVIRT [41] | 78.86 | 85.42 | 87.64 | 72.39 | 80.41 | 91.67 | 90.30 | 97.74 | 99.70 |
| GLoRIA [39] | 79.13 | 85.59 | 87.83 | 75.85 | 86.20 | 91.89 | 92.74 | 97.18 | 99.54 |
| BioViL [42] | 80.27 | 86.04 | 88.29 | 70.29 | 79.45 | 88.05 | 92.39 | 98.39 | 99.68 |
| MedKLIP [6] | 82.11 | 87.14 | 88.58 | 85.24 | 89.91 | 93.02 | 95.58 | 98.77 | 99.77 |
| MAVL [8] | 86.09 | 87.90 | 88.94 | 91.53 | 93.00 | 94.48 | 97.18 | 99.15 | 99.90 |
| Ours | 88.12 | 88.45 | 89.29 | 88.07 | 89.85 | 94.27 | 98.34 | 99.81 | 99.97 |

### v2

@@ Abstract
Recently, Multimodal Large Language Models (MLLMs) have demonstrated exceptional capabilities in visual understanding and reasoning across various vision-language tasks. However, we found that MLLMs cannot process effectively from fine-grained medical image data in the traditional Visual Question Answering (VQA) pipeline, as they do not exploit the captured features and available medical knowledge fully, results in MLLMs usually performing poorly in zero-shot medical disease recognition. Fortunately, this limitation does not indicate that MLLMs are fundamentally incapable of addressing fine-grained recognition tasks. From a feature representation perspective, MLLMs demonstrate considerable potential for tackling such challenging problems. Thus, to address this challenge, we propose LLaVA-RadZ, a simple yet effective framework for zero-shot medical disease recognition via utilizing the existing MLLM features. Specifically, we design an end-to-end training strategy, termed Decoding-Side Feature Alignment Training (DFAT) to take advantage of the characteristics of the MLLM decoder architecture and incorporate modality-specific tokens tailored for different modalities. Additionally, we introduce a Domain Knowledge Anchoring Module (DKAM) to exploit the intrinsic medical knowledge of large models, which mitigates the category semantic gap in image-text alignment. Extensive experiments demonstrate that our LLaVA-RadZ significantly outperforms traditional MLLMs in zero-shot disease recognition, achieving the comparable performance to the well-established and highly-optimized CLIP-based approaches.
@@ 1 Introduction
With the rapid advancement of deep learning technologies, an increasing number of studies have focused on their applications in medical disease diagnosis, yielding remarkable results (Chan et al., 2020; Jamshidi et al., 2020; Lee et al., 2022; Tran et al., 2021). However, these approaches typically rely on high-quality annotations provided by clinical experts. Unlike natural image datasets, annotating medical images is both costly and time-consuming. To address this challenge, recent research has explored methods based on paired medical images and textual reports, leveraging contrastive learning techniques. By minimizing the distance between paired samples while maximizing the distance between unpaired ones, these CLIP-based approaches enable zero-shot disease recognition, thereby reducing reliance on extensive medical data annotation to a certain extent. In our in-depth investigation of advanced zero-shot disease recognition methods in the medical domain, several representative CLIP-based models (Lai et al., 2024; Wu et al., 2023; Zhang et al., 2023b; Phan et al., 2024) have achieved significant performance improvements leveraging the capabilities of Large Language Models (LLMs) or incorporate expert domain knowledge to some extent, rather than fully leveraging the models’ intrinsic understanding capabilities.
@@
Recently, Multimodal Large Language Models (MLLMs) (Achiam et al., 2023; Team et al., 2023; Liu et al., 2023; Huang et al., 2024; 2025; You et al., 2025) have demonstrated remarkable capabilities across various user-oriented vision-language tasks, such as image comprehension and reasoning, offering new possibilities for zero-shot disease recognition in medical applications. Among these, LLaVA-Med (Li et al., 2024a) has exhibited exceptional domain-specific medical knowledge in dialogue-based tasks, indicating that it possesses a certain degree of medical expertise. However, a recent study (Zhang et al., 2024) found that MLLMs, i.e., LLaVA (Liu et al., 2023)) performed significantly worse than CLIP (Radford et al., 2021) on standard image classification tasks.
@@
To further validate this observation, we conducted zero-shot classification experiments using multiple MLLMs on five medical imaging datasets (see Tab. 1). The experimental results are consistent with previous findings, confirming that MLLMs exhibit suboptimal performance in image classification, particularly when dealing with complex medical images. To enhance the generalization capability of MLLMs in radiology disease recognition tasks, we employed a fine-tuning strategy and performed supervised fine-tuning on the MIMIC-CXR dataset (Johnson et al., 2019). Additionally, inspired by the work of (Zhang et al., 2024), we incorporated a series of optimizations. While these improvements yielded performance gains, the results remained inferior compared to CLIP-based models. This phenomenon raises a critical question: Can MLLMs effectively perform zero-shot disease recognition?
As shown in Fig. 1, we visualize the feature distributions of MAVL (Phan et al., 2024), LLaVA-Med (fine-tuned by the same dataset of our LLaVA-RadZ) and LLaVA-RadZ on the RSNA (Shih et al., 2019) dataset. The results indicate that MLLM exhibits strong feature extraction capabilities, comparable to the well-established MAVL in the domain. However, in the disease recognition task, MAVL significantly outperforms fine-tuned LLaVA-Med. We hypothesize that this performance gap arises because MLLMs fail to fully utilize the extracted features for effective disease identification via traditional VQA pipeline.
@@
Inspired by this, we propose a simple yet effective LLaVA-RadZ framework for zero-shot disease recognition using the MLLM features. Our proposed framework has the fundamental difference compared with previous CLIP-base methods and traditional MLLM VQA pipeline. As shown in Fig. 2, we design a dedicated MLLM feature-based framework to address zero-shot medical disease recognition. Our proposed framework effectively leverages pre-trained MLLM representations to overcome the inherent limitations of the traditional VQA pipeline on this task. Specifically, firstly, we introduce a new training strategy, Decoding-Side Feature Alignment Training (DFAT). Specifically, we introduce special tokens for both image and text modalities and leverage the autoregressive generation capability of the decoder architecture to extract global representations of images and texts. Additionally, we incorporate a cross-modal contrastive loss to optimize the model’s ability to learn discriminative features. Furthermore, to mitigate the semantic category gap encountered during fine-grained alignment between medical images and textual reports, we design a Domain Knowledge Anchoring Module (DKAM). DKAM utilizes the model’s intrinsic medical knowledge to extract the semantic information underlying disease categories, constructing disease description vectors that serve as an intermediary bridge to facilitate the alignment between medical images and textual reports, thereby establishing a stable relationship. To further enhance the correlation among medical images, textual reports, and disease categories, a category knowledge-guided loss strengthens the association between similar images and corresponding textual reports.
@@
We analyze the limitations of current MLLMs in addressing complex fine-grained medical disease recognition tasks, investigate the underlying causes of these constraints, and propose a novel end-to-end feature-based MLLM framework to mitigate these challenges. To the best of our knowledge, we are the first work in the field of medical disease recognition to explore how to use MLLM features directly to solve complex recognition problems.
We propose the tailored training strategy DFAT, and incorporate a cross-modal contrastive loss to optimize the model’s ability to achieve effective alignment between visual and textual features. Furthermore, we design a DKAM to leverage MLLM’s intrinsic medical knowledge and effectively mitigate semantic gap in image-text alignment, thereby enhancing category-level alignment.
@@ 2.1 Can Med-LLMs Be Good Medical Classifiers?
Previous studies have explored the classification capabilities of multimodal large language models (MLLMs), revealing that their performance on image classification tasks is often limited. For example, (Zhang et al., 2024) investigates the performance differences in classification between MLLMs and CLIP, focusing on factors such as inference strategies, training approaches, and datasets. Inspired by this work, we extend the exploration to zero-shot tasks in the medical domain. Unlike natural images and text, the relationship between medical images and reports is more complex. We seek to investigate whether large medical models, leveraging domain-specific knowledge, can achieve superior performance on medical zero-shot tasks.
We first evaluated two open-source MLLMs, i.e., LLaVA-1.5 (Liu et al., 2023) and LLaVA-Med (Li et al., 2024a), on five medical datasets in a zero-shot classification setting. The evaluation followed a general large-model classification approach, where the model selects the correct category from a set of candidate options. As shown in Tab. 1, these models demonstrated limited performance in disease classification tasks and failed to accurately identify various medical conditions. Given the potential knowledge limitations of these models, we further assessed the performance of more powerful proprietary MLLMs ( i.e., Qwen2.5-Max (Yang et al., 2024), Gemini-Pro (Team et al., 2023), and GPT-4o (Achiam et al., 2023)) on zero-shot medical disease recognition tasks. As shown in table 1, these models exhibited superior classification capabilities. However, they still lagged behind the state-of-the-art domain-specific methods in medical classification.
To enhance the generalization ability of MLLMs in radiology disease identification, we conducted Supervised Fine-Tuning (SFT) on LLaVA-1.5 (Liu et al., 2023) and LLaVA-Med (Li et al., 2024a) using the publicly available MIMIC-CXR dataset (Johnson et al., 2019). Surprisingly, the fine-tuned models did not achieve consistent performance improvements across the five datasets. In some cases, their classification performance even deteriorated. Further analysis of the model outputs revealed that MLLMs did not always focus on disease-specific information in radiology reports. Instead, they tended to overlearn the textual structures and linguistic patterns of the reports, which limited their classification capability. To mitigate this issue, we incorporated the Chain-of-Thought (CoT) prompting strategy and adjusted the model’s reasoning approach, inspired by the methodology of (Zhang et al., 2024), to optimize the model’s decision-making process. This approach led to moderate improvements in classification performance on medical datasets. Although the models have not yet reached optimal performance, the results suggest that MLLMs still hold significant potential for zero-shot medical disease recognition.
@@ 2.2 Motivation
To address this limitation, we propose the LLaVA-RadZ framework, introducing a novel end-to-end training strategy, Decoding-Side Feature Alignment Training (DFAT). This approach leverages the unique properties of the MLLM decoder architecture while incorporating modality-specific special tokens to facilitate effective interaction between medical images and textual features, ultimately achieving more robust cross-modal alignment. As illustrated in Fig. 1, we compare the feature distribution of our model with MAVL (Phan et al., 2024), the current state-of-the-art method, on the RSNA (Shih et al., 2019) dataset. The results clearly demonstrate that our model achieves better clustering of intra-class samples while enhancing inter-class separation, validating the effectiveness of our approach. Furthermore, we introduce the Domain Knowledge Anchor Module (DKAM), which harnesses the intrinsic medical knowledge of LLMs to bridge the semantic gap between images and text, enabling more precise disease classification.
@@ 2.3.2 Domain Knowledge Anchor Module
| Dataset | CheXpert | ChestXray-14 | COVIDx CXR-2 | RSNA Pneumonia | SIIM-ACR | |||||||||||
| MLLM | LLaVA-1.5 (7B) (Liu et al., 2023) | - | 7.50 | 8.28 | - | 3.33 | 6.92 | - | 53.14 | 50.28 | - | 40.53 | 55.34 | - | 23.66 | 50.36 |
| LLavA-Med (7B) (Li et al., 2024a) | - | 6.87 | 8.94 | - | 8.02 | 6.78 | - | 34.90 | 50.03 | - | 18.58 | 50.00 | - | 21.91 | 49.90 | |
| Qwen2.5-Max (Yang et al., 2024) | - | 32.23 | 67.97 | - | 19.04 | 76.19 | - | 75.91 | 76.81 | - | 43.58 | 43.59 | - | 64.70 | 72.57 | |
| Gemini-Pro (Team et al., 2023) | - | 35.01 | 76.08 | - | 14.16 | 77.78 | - | 62.84 | 62.90 | - | 44.23 | 51.43 | - | 61.43 | 72.03 | |
| GPT-4o (Achiam et al., 2023) | - | 45.85 | 81.14 | - | 19.85 | 81.55 | - | 50.93 | 77.08 | - | 54.20 | 65.33 | - | 64.57 | 72.11 | |
| LLavA-Med-7B + CoT (Zhang et al., 2024) | - | 8.90 | 26.23 | - | 8.33 | 20.46 | - | 27.12 | 26.55 | - | 49.59 | 43.80 | - | 54.06 | 51.07 | |
| LLavA-Med-7B + Inference (Zhang et al., 2024) | 71.00 | 44.85 | 75.45 | 64.30 | 21.73 | 70.86 | 71.07 | 69.84 | 60.39 | 77.51 | 69.85 | 72.90 | 71.25 | 68.26 | 71.27 | |
| Ours | LLaVA-RadZ | 73.36 | 48.59 | 82.15 | 72.61 | 27.91 | 84.64 | 84.36 | 77.53 | 74.58 | 86.98 | 76.18 | 83.28 | 89.92 | 79.57 | 84.38 |
| Method | ChestXray-14 | COVIDx CXR-2 | RSNA Pneumonia | SIIM-ACR | ||||||||
| AUC | F1 | ACC | AUC | F1 | ACC | AUC | F1 | ACC | AUC | F1 | ACC | |
| ConVIRT (Zhang et al., 2022) | 53.15 | 12.38 | 57.88 | 62.78 | 71.23 | 63.84 | 79.21 | 55.67 | 75.08 | 64.25 | 42.87 | 53.42 |
| GLoRIA (Huang et al., 2021) | 55.92 | 14.20 | 59.47 | 64.52 | 70.78 | 60.21 | 70.37 | 48.19 | 70.54 | 54.71 | 40.39 | 47.15 |
| BioViL (Boecking et al., 2022) | 57.82 | 15.64 | 61.33 | 61.40 | 70.92 | 58.20 | 84.12 | 54.59 | 74.43 | 70.28 | 46.45 | 68.22 |
| CheXzero (Tiu et al., 2022) | 66.99 | 21.99 | 65.38 | 73.13 | 76.13 | 71.45 | 85.13 | 61.49 | 78.34 | 84.60 | 65.97 | 77.34 |
| MedKLIP (Wu et al., 2023) | 72.33 | 24.18 | 79.40 | 76.28 | 76.54 | 71.96 | 86.57 | 63.28 | 79.97 | 89.79 | 72.73 | 83.99 |
| MAVL (Phan et al., 2024) | 73.50 | 26.25 | 82.77 | 83.86 | 81.73 | 78.07 | 86.91 | 63.41 | 82.42 | 92.04 | 77.95 | 87.14 |
| Ours | 72.61 | 27.91 | 84.64 | 84.36 | 77.53 | 74.58 | 86.98 | 76.18 | 83.28 | 89.92 | 79.57 | 84.38 |
| Method | RSNA Pneumonia | Pneumothorax | COVIDx CXR-2 | ||||||
| 1% | 10% | 100% | 1% | 10% | 100% | 1% | 10% | 100% | |
@@
| ConVIRT (Zhang et al., 2022) | 78.86 | 85.42 | 87.64 | 72.39 | 80.41 | 91.67 | 90.30 | 97.74 | 99.70 |
| GLoRIA (Huang et al., 2021) | 79.13 | 85.59 | 87.83 | 75.85 | 86.20 | 91.89 | 92.74 | 97.18 | 99.54 |
| BioViL (Boecking et al., 2022) | 80.27 | 86.04 | 88.29 | 70.29 | 79.45 | 88.05 | 92.39 | 98.39 | 99.68 |
| MedKLIP (Wu et al., 2023) | 82.11 | 87.14 | 88.58 | 85.24 | 89.91 | 93.02 | 95.58 | 98.77 | 99.77 |
| MAVL (Phan et al., 2024) | 86.09 | 87.90 | 88.94 | 91.53 | 93.00 | 94.48 | 97.18 | 99.15 | 99.90 |
| Ours | 88.23 | 88.57 | 89.49 | 88.42 | 89.96 | 94.50 | 98.32 | 99.80 | 99.96 |